Find a pair of polar coordinates of the point whose cartesian coordinates are (3, -3 sqrt (3) )
Answer by stanbon(75887)

Find a pair of polar coordinates of the point whose cartesian coordinates are (3, -3 sqrt (3) )
----
Notice: The point is in QIV
----------------------------------
r = sqrt[3^2 + (3sqrt(3))^2] = sqrt[9+27] = sqrt(36) = 6
-----
theta = arctan[-3sqrt(3)/3] = arctan(-sqrt(3) = (11/6)pi
------------
Cheers,
Stan H.
-------------